# Retina Vision Transformer (RetinaViT): Introducing Scaled Patches into Vision Transformers

### v1

@@ 4 Experimental Evaluation
    The positional embedding calculations is highly reusable, as the embedding process is independent of the content of the images.
@@
    The scaling of the input image should be performed in the processing pipeline, instead of repeatedly inside the model.
The authors have not implemented these optimisations for a cleaner comparison with the original ViT implementation.
@@ 5 Discussion
The architectural changes in RetinaViT have a few theoretical implications.
5.1 Conceptual Expansion of Input Dimensionality
With the attention mechanism, Transformer effectively transforms an input sentence into a parse tree [5]. In recurrent models such as Long Short-term Memory (LSTM) [13], tokens in the input are fed into the model sequentially by their order of appearance in the sentence. But with attention, it is possible for a token with a high attention score to be effectively processed before preceding tokens. This is equivalent to allowing Transformers to move freely on a parse tree of the input sentence, and process tokens in an arbitrary order.
Similarly, by introducing scaled patches into ViT, RetinaViT transforms the input image, which is effectively the bottom layer of an image pyramid, into the full image pyramid, and allows attention to decide the order in which to process the patches. This is illustrated in fig. 6.
@@ 5.1 Conceptual Expansion of Input Dimensionality
(b) A flattened image pyramid with a potential attention path.
@@
This effectively expands the dimensionality of the input from two to three, as illustrated in fig. 4. Notice this expansion is conceptual only, since the format of the input has not changed, and there is no new information added to the input vector. The input to the Transformer Encoder is still a sequence of embedded patches, and all information in the patches is derived from the original image.
@@ 5.2.1 Scale-invariant Features
Traditionally, computer vision models, such as Scale-invariant Feature Transform (SIFT) [17], look for scale-invariant features that are preserved when the scale of the image is changed. This allows the recognition of the same features, and subsequently objects, in a close up as well as a further away image.
CNN is not innately scale-invariant [30]. Although it may be able to capture scale-invariant features to some extent, this is more due to the fact that each convolution layer has multiple kernels, and some of them may capture the same scale-invariant features at different scales. The original ViT is not scale-invariant either, since the patches are all from an image of one single scale.
RetinaViT is closer to CNN than ViT, because its input contains the same image at multiple different scales. This means that each scale-invariant feature is guaranteed to be seen by RetinaViT $n$ times, where $n$ is the number of layers in the input image pyramid. As a result, RetinaViT is stronger than both CNN and ViT in recognizing scale-invariant features, and has a higher chance of detecting the same objects at various scales.
@@ 5.3 On Inductive biases
The authors of the original paper on ViT state that fewer inductive biases are present in the ViT setup, and explain the improved performance of ViT over CNN with “large scale training trumps inductive bias” [9].
@@
Notice that RetinaViT has not hand-crafted any attention patterns, since how attention scores are assigned to each patch is still learned during the training process. Therefore no extra inductive bias is introduced on this front.
One inductive bias that is indeed reintroduced into RetinaViT, in a very subtle way, is that as the network goes deeper, only the more important features are forwarded to the next layer.
In CNN, as the network goes deeper, the size of the input to each layer shrinks, whereas the receptive field size of each convolution kernel increases. In the final layer, only the most important features are preserved after the Max Pooling processes throughout network.
In some other models, such as the spiking neural network framework Assembly Calculus [19], there is a Random Projection and Cap primitive, where one group of neurons randomly connect to downstream neurons, and only the connections that are the most consistent with Hebbian plasticity are preserved. This is also a selection process, with a cap on information capacity.
In RetinaViT, patches from multiple scales are added to the input of the Transformer Encoder, so there is more information in the first layer than the original ViT. But since the size of the Transformer blocks has not changed, there is effectively a capping process, where the most important information from patches from all five scales are selected, and forwarded to deeper layers.
6 Future Work
The authors have formulated several hypotheses based on RetinaViT, and plan to investigate these in future work.
6.1 Emergence of Vertical Pathways
Learning part-whole relationships, or how features extracted from fine-grained details such as texture can be bound with those from higher level characteristics such as contour, has always been a focus of studies. For example, CapsNet [22] uses agreement between activity vectors of capsules, and GLOM [12] uses spatially local columns, to bind features from different scales to a single processing pathway.
In CNN, due to the preservation of the spatial arrangement of features across layers, binding of low and high level features is achieved by design. In ViT, because of the flexibility in processing order provided by the attention mechanism, there is not such a binding mechanism.
In RetinaViT, there is no explicit binding mechanism either, but the authors hypothesize that including patches from different scales can enable the emergence of different vertical pathways inside RetinaViT. It is possible for RetinaViT to form a pathway where the initial information comes from patches containing low spatial frequency components, and only interact with another pathway that initiates from patches containing high spatial frequency components deeper down the network. The authors plan to look for such pathways in future studies.
6.2 Attention Patterns
CNNs have a less-discussed inductive bias, that processing starts from more local details, and ends at the higher level features and characteristics. This is contrary to how humans process visual input. As discussed in [23], humans process low spatial frequency components quicker than high spatial frequency components. This helps perceive the “gist” of the scene as a whole first, before turning to more nuanced matters.
With RetinaViT, since low and high spatial frequency components are fed into the Transformer Encoder at the same time, it is possible for low spatial frequency components to be attended to, before high spatial frequency components are. The authors plan to investigate such patterns in future work as well.
@@ 7 Conclusion
In this paper, we investigated the hypothesis that providing information from the same image in multiple scales to a neural network may allow it to learn a better model for a visual task. To test this idea in a controlled way, we devised an approach whereby we supply all patches from a hierarchy of scaled images, referred to as an image pyramid, as input to the standard Vision Transformer model. We showed experimentally that this approach can improve performance on a benchmark dataset. Several directions for future work arise from the discussion of our approach, which we plan to investigate further.

### v2

@@ 4.1 Magnitude of Values
    In the initial layer, low spatial frequency components, or low resolution patches, received higher attention weights. In other words, low spatial frequency components are more important in the first layer. In future layers, the attention weights assigned to high spatial frequency components, or high resolution patches, increased significantly.
@@
    This trend is more apparent in the magnitude of the attention scores. In the first layer, only low spatial frequency components received high attention scores, but from the second layer on, layer attention shifted to high spatial frequency components, which overtook the low spatial frequency components as the dominant patches in the layers.
  3. 3.
    The sum of attention scores and the original inputs can be interpreted as a reflection of the changes caused by the attention scores on the original inputs. It is apparent from the experiment results that low spatial frequency components are more important in the input to the MLP.
Figure 6: Results averaged from probing the trained models for magnitude of attention weights (left), attention scores (centre), and the sum of the attention scores with the inputs from the residual connection (right). Results are ordered vertically from the input layer (at top) to layer 11 (at bottom). Red markers on the horizontal axes denote boundaries between patches of different resolutions.
With these observations, shown in Figure 6, we conclude that low spatial frequency components play a significant role in the early layers of RetinaViT, similar to the fact that low spatial frequency components are processed earlier than high spatial frequency components in the human visual system. It is also worth noting that this preference for low spatial frequency components is not an inductive bias imposed on the model, but one that emerged by itself in the training process.
We hypothesise that this preference can be explained by the fact that understanding the gist of the scene, which is captured better in low spatial frequency components, informs the interpretation of high spatial frequency components.
4.2 Robustness to Size Reduction
We compared the robustness of RetinaViT against reduction in the number of layers compared to the original ViT. We reduced the number of layers gradually from the initial configuration, with 12 layers, down to only 2 layers, and observed the performance changes of the two models, shown in Table 1.
Model Number of Layers in the Model
2 4 6 8 10 12
ViT 46.4% 64.5% 71.1% 73.7% 75.4% 76.5%
RetinaViT 47.2% 65.1% 71.2% 73.8% 75.5% 76.7%
Difference +0.8% +0.6% +0.1% +0.1% +0.1% +0.2%
Table 1: Top-1 prediction accuracies of the original ViT and RetinaViT on the val validation set, with different numbers of layers. Each configuration is executed once.
We see from Table 1 that whilst RetinaViT did not provide a huge improvement in absolute prediction accuracy at the full configuration, it did improve the robustness of the model when the number of layers is reduced. We hypothesise that RetinaViT achieved this by directly capturing the low spatial frequency components, or high level, structural information of the scene.
@@ 5 Discussion
5.1 Emergence of Low to High Processing Order
With the attention mechanism, Transformer effectively transforms an input sentence into a parse tree Bird et al. (2009). In recurrent models such as Long Short-term Memory (LSTM) Hochreiter & Schmidhuber (1997), tokens in the input are fed into the model sequentially by their order of appearance in the sentence. But with attention, it is possible for a token with a high attention score to be effectively processed before preceding tokens. This is equivalent to allowing Transformers to move freely on a parse tree of the input sentence, and process tokens in an arbitrary order.
Similarly, by introducing scaled patches into ViT, RetinaViT transforms the input image, which is effectively the bottom layer of an image pyramid, into the full image pyramid, and allows attention to decide the order in which to process the patches. This is illustrated in Figure 7.
@@ 5.1 Emergence of Low to High Processing Order
(b) A flattened image pyramid with a hypothetical attention path.
@@
Since the preference for low spatial frequency components in early layers, and that for high spatial frequency components in later layers emerged by itself, we hypothesise that RetinaViT learned to process a visual parse tree from the structural nodes to the terminal nodes, in reverse of the processing order of CNNs.
@@ 5.2.1 Scale-invariant Features
Traditional computer vision models, such as Scale-invariant Feature Transform (SIFT) Lowe (1999), look for scale-invariant features that are preserved when the scale of the image is changed. This allows the recognition of the same features, and subsequently objects, in a close up as well as a further away image.
CNN is not innately scale-invariant Xu et al. (2014). Although it may be able to capture scale-invariant features to some extent, this is more due to the fact that each convolution layer has multiple kernels, and some of them may capture the same scale-invariant features at different scales. The original ViT is not scale-invariant either, since the patches are all from an image of one single scale.
RetinaViT is closer to CNN than ViT in this front, because its input contains the same image at multiple different scales, meaning that each scale-invariant feature is guaranteed to be seen by RetinaViT $n$ times, where $n$ is the number of layers in the input image pyramid. As a result, RetinaViT is stronger than both CNN and ViT in recognising scale-invariant features, and has a higher chance of detecting the same objects at various scales.
@@ 5.3 On Inductive biases
The authors of the original paper on ViT state that fewer inductive biases are present in the ViT setup, and explain the improved performance of ViT over CNN with “large scale training trumps inductive bias” Dosovitskiy et al. (2021).
@@
Notice that RetinaViT has not introduced any hand-crafted attention patterns, in other words inductive biases, into the model, since how attention scores are assigned to each patch is still learned during the training process.
One inductive bias that is indeed reintroduced into RetinaViT, in a very subtle way, is that as the network goes deeper, only the more important features are forwarded to the next layer. This inductive bias is present in many models. In CNNs, only the most important features are preserved after the Max Pooling processes throughout network. In the spiking neural network framework Assembly Calculus Papadimitriou et al. (2019), there is a Random Projection and Cap primitive, where a group of neurons randomly connect to downstream neurons, and only the connections that are the most consistent with Hebbian plasticity are preserved. This is a selection process, with a cap on information capacity. In RetinaViT, patches from multiple scales are added to the input of the Transformer Encoder, but since the size of the Transformer blocks has not changed, there is effectively a capping process, where the most important information from patches from all five scales are selected, and forwarded to deeper layers.
@@ 6 Conclusion
In this paper, we investigated the effect of supplying all patches from a hierarchy of scaled images to Vision Transformers, and showed that the resulting model, RetinaViT, attends more to low spatial frequency components, or low resolution patches, in the initial layer, and then switches to the high spatial frequency components in subsequent layers. This processing order emerged by itself without any imposed inductive bias, and is consistent with how the human visual system processes a scene.
We argue that given that the low to high frequency processing order emerged by itself in a model with minimal inductive bias regarding processing orders, model architectures where this processing order is reinforced as inductive biases have strong potential, and are thus worth future exploration.
7 Acknowledgement
This research was undertaken with the assistance of resources from the National Computational Infrastructure (NCI Australia), an NCRIS enabled capability supported by the Australian Government.